OpenStudy (anonymous):

7-1) Reduce the following rational expression to lowest terms, if possible. 16a+48 _______ 4a+12

OpenStudy (anonymous):

\[\frac{16a+48}{4a+12}=\frac{4(4a+12)}{4a+12}=4\]